OpenStudy (anonymous):

if ||u||=1 and ||v||=5, and the vectors make angle of 130 and 180 respectively, find the component form of the sum of u and v.

OpenStudy (tkhunny):

I'm going to guess that is has something to do with \(\cos(130º) + 5\cos(180º)\).

OpenStudy (anonymous):

That is indeed the x to one of the possible solutions... though i doubt it is the answer.

OpenStudy (tkhunny):

Okay, can you write 'y' in a similar fashion?

OpenStudy (anonymous):

None that i can identify

OpenStudy (anonymous):

Try replacing cosine with sine for the vertical component.

OpenStudy (anonymous):

Was it seriously that easy... that is an answer...

OpenStudy (anonymous):

Dear god i over thought that too much.

OpenStudy (tkhunny):

Who cares if it is AN answer Is it correct?

OpenStudy (tkhunny):

You could figure out that \(\sin(180º) = 0\) and just skip that part.